# InSpire: Vision-Language-Action Models with Intrinsic Spatial Reasoning

### v2

@@ Abstract
Leveraging pretrained Vision-Language Models (VLMs) to map language instruction and visual observations to raw low-level actions, Vision-Language-Action models (VLAs) hold great promise for achieving general-purpose robotic systems. Despite their advancements, existing VLAs tend to spuriously correlate task-irrelevant visual features with actions, limiting their generalization capacity beyond the training data. To tackle this challenge, we propose Intrinsic Spatial Reasoning (InSpire), a simple yet effective approach that mitigates the adverse effects of spurious correlations by boosting the spatial reasoning ability of VLAs. Specifically, InSpire redirects the VLA’s attention to task-relevant factors by prepending the question “$\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot?}$” to the language instruction and aligning the answer “$\mathsf{right/left/up/down/front/back/grasped}$” and predicted actions with the ground-truth. Notably, InSpire can be used as a plugin to enhance existing autoregressive VLAs, requiring no extra training data or interaction with other large models. Extensive experimental results in both simulation and real-world environments demonstrate the effectiveness and flexibility of InSpire.
1 Introduction
@@
In recent years, Vision-Language Models (VLMs) [44, 37] have demonstrated remarkable capabilities across a diverse set of tasks, including image captioning and visual question answering (VQA). These advances have paved the way for Vision-Language-Action (VLA) models, which utilize pretrained VLMs to directly map natural language commands and visual inputs to low-level motor actions, offering a promising pathway toward general-purpose robotic systems [5, 15, 26, 32].
@@
Despite their advancements, state-of-the-art VLAs tend to spuriously correlate task-irrelevant visual features with actions, overlooking vital elements such as language instructions and spatial relations in visual observations, which hinders their ability to generalize beyond the training data distribution. As shown in Fig. 1 (a), the VLA—trained via direct observation-to-action mapping—fails to accurately identify task-specific objects and model their spatial relationships with the robot. Instead, it allocates more attention to task-irrelevant regions when predicting actions. Understanding and reasoning about spatial relationships are crucial skills that empower humans to solve complex tasks: if asked the same question, they would first try to assess the directions (or coarse locations) of the “black bowl” and the “plate” relative to their hands. This underscores the importance of leveraging spatial reasoning as a bridge to capture the causal relationships between observations and actions, thereby allowing VLAs to produce more accurate and robust robot actions, as illustrated in Fig. 1 (b). Prior efforts generally resort to auxiliary training data [48] or other large models [41, 28] to enhance chain-of-thought (or step-by-step) reasoning capabilities. While these approaches offer partial improvements in spatial reasoning, they often suffer from limitations in efficiency and generalizability. Therefore, a research question remains open in the field of robot learning: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] Without relying on extra data or interacting with other large models, can we enhance the spatial reasoning capabilities of VLAs to resolve spurious correlations?
@@
Figure 1: (a) VLAs typically predict actions relying on Spurious Correlations learned by the direct observation-to-action mapping mechanism. (b) The core idea of our InSpire method that tackles spurious correlations by boosting the spatial reasoning capabilities of VLAs. (c) InSpire can be used as a plugin to improve state-of-the-art VLAs on both seen and unseen tasks across simulation and real-world environments. For illustrative purposes, we omit the tokens of the language instruction “pick up the black bowl next to the plate and place it on the plate”.
@@
Our solution to the question is Intrinsic Spatial Reasoning (InSpire), a simple yet effective approach that boosts VLAs’ spatial reasoning capabilities to mitigate the adverse effects of spurious correlations on their generalization in novel scenarios. As presented in Fig. 2, our InSpire approach redirects the model’s attention from spurious factors to task-relevant ones by simply appending the question “$\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot?}$” before the language instruction and aligning the VLA’s generated answer “$\mathsf{right/left/up/down/front/back/grasped}$” and predicted actions with the ground-truth. By using this spatial reasoning VQA task as the bridging “language” between observations and actions, InSpire equips VLAs with the ability to understand and reason about spatial relationships without the need to collect auxiliary training data or interact with other large models. Notably, the InSpire approach is fully compatible with existing autoregressive VLAs and can be seamlessly integrated as a plugin to enhance their performance.
Effectiveness and Flexibility. We conduct extensive evaluations in both simulation and real-world environments, on top of two state-of-the-art VLAs, including miniVLA-VQ[3] (a lightweight version of OpenVLA[15]) and $\pi_{0}$-FAST [32] (an upgraded version of $\pi_{0}$[4]). The achieved results demonstrate InSpire’s effectiveness and flexibility. Remarkably, InSpire enhances the absolute success rate of miniVLA-VQ by 6.2% on seen tasks and 10% on unseen tasks in the simulation environment, and improves $\pi_{0}$-FAST by 25% on seen tasks and 26% on unseen tasks in the real-world environment (see Fig. 1 (c)). Moreover, building upon the miniVLA-VQ architecture, we pretrain a 1B-parameter VLA named InspireVLA-1B on the LIBERO [21] benchmark, showcasing superior performance and computational efficiency compared to state-of-the-art reasoning-based VLAs. Code, pretrained models and demos are publicly available [^1] https://github.com/InspireVLA/Inspire.
@@
Figure 2: Overview of our InSpire approach. InSpire boosts the VLA’s spatial reasoning ability by appending the question “$\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot?}$” before the language instruction and aligning the VLA’s answer “$\mathsf{right/left/up/down/front/back/grasped}$” and predicted actions with the ground-truth. InSpire is compatible with existing autoregressive VLAs.
2 Proposed Approach
2.1 Preliminaries and Problem Statement
@@
2.2 Intrinsic Spatial Reasoning
We hypothesize that models resort to spurious correlations when task-relevant factors are not explicit in the input. Such latent factors are difficult to learn, as neural networks preferentially learn simpler patterns first [1]. For instance, in a task such as “put the bowl in the drawer of the cabinet” (see Figure 2), the small size of the bowl might lead the model to infer actions from irrelevant distractors or background elements rather than the object of interest. The central insight behind our method is to enable the model to first extract salient, task-relevant information from observations—a simpler pattern to discern—and then utilize this information as an additional input for action generation.
Method Overview. Formally, we denote $u^{\prime}=f(u)$ as an extracted representation that processes and summarizes high-level information from the task-relevant factors $u$. The goal is to learn two policies: an extraction policy $\pi_{u^{\prime}}:\mathcal{O}\times\mathcal{L}\rightarrow\mathcal{U}^{\prime}$ which maps observation and task instruction to the extracted task-relevant representations, and an action policy $\pi_{\theta}:\mathcal{O}\times\mathcal{L}\times\mathcal{U}^{\prime}\rightarrow\mathcal{A}$ which maps observation, task instruction, and the extracted task-relevant representations to actions. Consequently, the model outputs actions via a two-step process instead of a single step:
@@ 2.2 Intrinsic Spatial Reasoning
Modeling Task-relevant Factors via Spatial Reasoning VQA. To establish a reliable representation $u^{\prime}$ of task-relevant factors, we leverage the extensive, text-based world knowledge inherent in VLMs—the models from which contemporary VLAs are often pretrained. Exploiting text as this representation $u^{\prime}$, we propose Intrinsic Spatial Reasoning (InSpire) that performs explicit spatial reasoning regarding the robot and objects of interest prior to the inference of subsequent actions.
As depicted in Figure 2, we employ a VLA that dually functions as an extraction policy $\pi_{u^{\prime}}$ and an action policy $\pi_{\theta}$. Given an initial language instruction $l$, the InSpire framework first introduces a textual question $q$ to probe the spatial relationships between objects mentioned in $l$ and the robot, e.g., “$\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot}?$”. Object names are identified within $l$ using the natural language toolkit [24]. The VLA, operating as the extraction policy $\pi_{u^{\prime}}$, takes $q$, the current visual observation $o$, and the language instruction $l$ as input to produce a textual answer $g$, which is constrained to a predefined set of valid options of coarse-grained directions, as detailed in Figure 3. The combination of the self-generated question and its corresponding answer constitutes an extracted textual representation $u^{\prime}=[q,g]$. This representation $u^{\prime}$ is subsequently passed to the same VLA, now serving as the action policy $\pi_{\theta}$, which outputs the final robot actions. Because this spatial reasoning VQA task is specifically formulated to correlate highly with objects of interest and the requisite actions, the resulting textual pair $u^{\prime}$ provides a distilled, abstract representation of task-relevant factors, simplifies the learning challenge for the action policy, and mitigates the learning of spurious correlations.
@@
Automated Rule-based Object Direction Labeling. To facilitate model training, ground-truth spatial relationships are integrated into the training datasets. Each training datapoint is thus augmented to $(o_{i},a_{i},u^{\prime}_{i})$, where $u^{\prime}_{i}$ encompasses self-generated questions and their corresponding ground-truth spatial relationship answers. These ground-truth relationships are derived by first determining the positions of the robot’s end-effector/gripper and relevant objects, then inferring their relative spatial arrangement, as illustrated in Figure 3. In simulations, object locations are readily available. In real-world settings, the end-effector’s position at the instant of gripper closure/opening is recorded and used as a proxy for the object’s position when establishing these relationships (see Appendix B). Given an end-effector position $[x_{i},y_{i},z_{i}]$ and an object position $[x_{0},y_{0},z_{0}]$, the position difference is calculated as $\bm{d}=[x_{i}-x_{0},y_{i}-y_{0},z_{i}-z_{0}]$. The coarse-grained spatial relationship is then determined by the axis corresponding to the component of $\bm{d}$ with the largest absolute value. For instance, if $|x_{i}-x_{0}|$ is the maximum component magnitude in $\bm{d}$, the relationship is classified as “left” or “right” based on the sign of $x_{i}-x_{0}$. When the object is grasped, as indicated by gripper closure, it is directly labeled with the text “grasped”. This rule-based strategy, applied from a third-person viewpoint, generates the ground-truth labels used for supervising the VLA’s spatial reasoning. During training, the autoregressive loss is also applied to tokens corresponding to spatial reasoning answers, alongside the primary action tokens; this additional loss encourages the model to predict the correct textual descriptions of these spatial relationships. Our experiments show that this relatively coarse-grained spatial reasoning is nonetheless effective for guiding the action generation process (see Section 3.4).
3 Experiments
In this section, we conduct extensive experiments to answer the following questions: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] 1) Can InSpire enhance VLAs on tasks from both simulation and real-world environments?
@@ 3 Experiments
We answer the first question in Sections 3.1 and 3.2, the second question in Section 3.3, the third question in Section 3.4, and the fourth question in Section 3.5. More details about the baseline VLAs, evaluation tasks and hyperparameters used in our simulation and real-world experiments are presented in Appendix C. Additional experimental results are provided in Appendix D.
3.1 Simulation Experiments
We perform simulation experiments using the LIBERO [21] environment/benchmark. Concretely, we use manipulation tasks from the five datasets—LIBERO-90, LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long/10—which feature a diverse range of objects, scene layouts, and language instructions. The VLA model is jointly trained on the 90 tasks in the LIBERO-90 dataset, with each task consisting of 50 demonstrations. To comprehensively assess InSpire’s capabilities, we evaluate the model on both seen tasks from LIBERO-90 and unseen tasks from the other four datasets. Unlike prior works [48, 41] that artificially create unseen tasks for each dataset in the simulation environment, we sample novel tasks from publicly available out-of-distribution datasets, ensuring result reproducibility and fair comparisons.
Experimental Setup. We perform evaluations using two state-of-the-art models: miniVLA-VQ[3] (a lightweight version of OpenVLA [15] with 1B parameters) and $\pi_{0}$-FAST [32] (an upgrated version of $\pi_{0}$ [4] with 3B parameters). In our experiments, the miniVLA-VQ model is pretrained from scratch, whereas $\pi_{0}$-FAST undergoes full-finetuning on the LIBERO-90 dataset. To ensure fair comparisons, both models adopt the same hyperparameters (e.g., learning rate and training step) as those employed for our InSpire. In addition to comparing InSpire with the two baseline VLAs, we evaluate it against the 7B-parameter OpenVLA model [15], the vanilla miniVLA and its two other variants miniVLA-VQ-history and miniVLA-VQ-wrist presented in the miniVLA project homepage  [^2] https://ai.stanford.edu/blog/miniVLA. Since the LIBERO-90 pretrained OpenVLA model has not been released, we directly borrow the reported results on seen tasks from the project homepage. We train the model with 3 random seeds, and conduct 100 trials per task during inference.
Results and Analysis. We have several key observations from the results in Table 1. 1) Our InSpire achieves the best or comparable performance compared to the two strong baseline VLAs on both seen and unseen tasks. Remarkably, InSpire enhances the absolute success rates of miniVLA-VQ by 6.2% and 10% on seen and unseen tasks, respectively. This underscores InSpire’s effectiveness and flexibility in tackling the adverse effects of the spurious correlation issue in VLAs, thereby enabling them to produce more accurate and robust robot actions. 2) The established performance gains of InSpire on $\pi_{0}$-FAST are not as significant as that on miniVLA-VQ. The primary reason could be that we train miniVLA-VQ from scratch but finetune the pretrained $\pi_{0}$-FAST. With substantially less data than $\pi_{0}$-FAST’s pretraining dataset, LIBERO-90 limits the ability of InSpire to enhance $\pi_{0}$-FAST’s spatial reasoning capability. 3) InSpire yields significantly better performance on LIBERO-Spatial and LIBERO-Object tasks compared to LIBERO-Goal and LIBERO-Long tasks, demonstrating its strength in handling spatial relationships and object interactions. The results on LIBERO-Long tasks also reveal InSpire’s inability to improve VLAs’ high-level task planning capacity. 4) SpatialVLA, as well as other state-of-the-art VLAs, exhibit poor generalization performance on tasks from the four unseen datasets, indicating that these models fail to capture true causal relationships between observations and actions, leading to poor generalization beyond the distribution of their training data.
Table 1: LIBERO Performance. Success rates (%) of the state-of-the-art VLAs miniVLA-VQ [3] and $\pi_{0}$-FAST [32] integrated w/ or w/o InSpire on the LIBERO benchmark. models borrowed from [3]. our models pretrained from scratch on LIBERO-90. our models finetuned on LIBERO-90.

### v3

@@ Abstract
Leveraging pretrained Vision-Language Models (VLMs) to map language instruction and visual observations to raw low-level actions, Vision-Language-Action models (VLAs) hold great promise for achieving general-purpose robotic systems. Despite their advancements, existing VLAs tend to spuriously correlate task-irrelevant visual features with actions, limiting their generalization capacity beyond the training data. To tackle this challenge, we propose Intrinsic Spatial Reasoning (InSpire), a simple yet effective approach that mitigates the adverse effects of spurious correlations by boosting the spatial reasoning ability of VLAs. Specifically, InSpire redirects the VLA’s attention to task-relevant factors by prepending the question $\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot?}$” to the language instruction and aligning the model’s output answer $\mathsf{right/left/up/down/front/back/grasped}$” and predicted actions with ground-truth. Notably, InSpire can be used as a plugin to enhance existing autoregressive VLAs, requiring no extra training data or interaction with other large models. Extensive experimental results in both simulation and real-world environments demonstrate the effectiveness and flexibility of InSpire.
I Introduction
@@
In recent years, Vision-Language Models (VLMs) [41, 35] have demonstrated remarkable abilities across a diverse set of tasks, including image captioning and visual question answering (VQA). These advances have paved the way for Vision-Language-Action (VLA) models, which use pretrained VLMs to directly map language commands and visual inputs to low-level motor actions, offering a promising pathway toward general-purpose robotic systems [5, 17, 26].
@@
Despite their advancements, state-of-the-art VLAs tend to spuriously correlate task-irrelevant visual features with actions, overlooking vital elements such as language instructions and spatial relations in visual observations, which hinders their ability to generalize beyond the training data distribution. As shown in Fig. 1 (a), the VLA—trained via direct observation-to-action mapping—fails to accurately identify task-specific objects and model their spatial relationships with the robot. Instead, it allocates more attention to task-irrelevant regions when predicting actions. Understanding and reasoning about spatial relationships are crucial skills that empower humans to solve complex tasks: if asked the same question, they would first try to assess the directions (or coarse locations) of the “black bowl” and the “plate” relative to their hands. This underscores the importance of leveraging spatial reasoning as a bridge to capture the causal relationships between observations and actions, thereby allowing VLAs to produce more accurate and robust robot actions, as illustrated in Fig. 1 (b). Prior efforts generally resort to auxiliary training data [45] or other large models [38, 28] to enhance chain-of-thought (or step-by-step) reasoning capabilities. While these approaches offer partial improvements in spatial reasoning, they often suffer from limitations in efficiency and generalizability. We thus raise the following question: Without relying on extra data or interacting with other large models, can we enhance the spatial reasoning capabilities of VLAs to resolve spurious correlations?
@@
Figure 1: (a) VLAs typically predict actions relying on Spurious Correlations learned by the direct observation-to-action mapping mechanism. (b) The core idea of our proposed InSpire method that tackles spurious correlations by boosting the spatial reasoning capabilities of VLAs.
Figure 2: Overview of our InSpire approach. InSpire boosts the VLA’s spatial reasoning ability by appending the question $\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot?}$” before the language instruction and aligning the VLA’s answer $\mathsf{right/left/up/down/front/back/grasped}$” and predicted actions with the ground-truth.
@@
Our solution the question is Intrinsic Spatial Reasoning (InSpire), a simple yet effective approach that boosts VLAs’ spatial reasoning capabilities to mitigate the adverse effects of spurious correlations on their generalization in novel scenarios. As presented in Fig. 2, our proposed InSpire approach redirects the model’s attention from spurious factors to task-relevant ones by simply appending the question $\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot?}$” before the language instruction and aligning the VLA’s generated answer $\mathsf{right/left/up/down/front/back/grasped}$” and predicted actions with the ground-truth. By using this spatial reasoning VQA task as the bridging “language” between observations and actions, InSpire equips VLAs with the ability to understand and reason about spatial relationships without the need to collect auxiliary training data or interact with other large models. Notably, the InSpire approach is fully compatible with existing autoregressive VLAs and can be seamlessly integrated as a plugin to enhance their performance. We conduct extensive evaluations using both simulation benchmarks (LIBERO [22] and CALVIN [27]) and real-world tasks, the achieved results on two state-of-the-art VLAs—miniVLA-VQ[3] (a lightweight version of OpenVLA[17]) and $\pi_{0}$-FAST [32] (an upgraded version of $\pi_{0}$[4]), prove InSpire’s effectiveness and flexibility.
@@
II Proposed Approach
II-A Preliminaries and Problem Statement
@@
II-B Intrinsic Spatial Reasoning
We hypothesize that models resort to spurious correlations when task-relevant factors are not explicit in the input. Such latent factors are difficult to learn, as neural networks preferentially learn simpler patterns first [1]. For instance, in a task such as “put the bowl in the drawer of the cabinet” (see Fig. 2), the small size of the bowl might lead the model to infer actions from irrelevant distractors or background elements rather than the object of interest. The central insight behind our method is to enable the model to first extract salient, task-relevant information from observations—a simpler pattern to discern—and then utilize this information as an additional input for action generation.
II-B1 Method Overview
Formally, we denote $u^{\prime}=f(u)$ as an extracted representation that processes and summarizes high-level information from the task-relevant factors $u$. The goal is to learn two policies: an extraction policy $\pi_{u^{\prime}}:\mathcal{O}\times\mathcal{L}\rightarrow\mathcal{U}^{\prime}$ which maps observation and task instruction to the extracted task-relevant representations, and an action policy $\pi_{\theta}:\mathcal{O}\times\mathcal{L}\times\mathcal{U}^{\prime}\rightarrow\mathcal{A}$ which maps observation, task instruction, and the extracted task-relevant representations to actions. Consequently, the model outputs actions via a two-step process instead of a single step:
@@ II-B1 Method Overview
II-B2 Modeling Task-relevant Factors via Spatial Reasoning VQA
To establish a reliable representation $u^{\prime}$ of task-relevant factors, we use the extensive, text-based world knowledge inherent in VLMs—the models from which contemporary VLAs are pretrained. Exploiting text as this representation $u^{\prime}$, we propose Intrinsic Spatial Reasoning (InSpire) that performs explicit spatial reasoning regarding the robot and objects of interest prior to the inference of subsequent actions.
As shown in Fig. 2, we employ a VLA that dually functions as an extraction policy $\pi_{u^{\prime}}$ and an action policy $\pi_{\theta}$. Given an initial language instruction $l$, InSpire first introduces a textual question $q$ to probe the spatial relationships between objects mentioned in $l$ and the robot, e.g., $\mathsf{In\,\,which\,\,direction\,\,is\,\,the\,\,[object]\,\,relative\,\,to\,\,the\,\,robot}?$”. Object names are identified within $l$ using the natural language toolkit [24]. The VLA, operating as the extraction policy $\pi_{u^{\prime}}$, takes $q$, the current visual observation $o$, and the language instruction $l$ as input to produce a textual answer $g$, which is constrained to a predefined set of valid options of coarse-grained directions, as detailed in Fig. 3. The combination of the self-generated question and its corresponding answer constitutes an extracted textual representation $u^{\prime}=[q,g]$. This representation $u^{\prime}$ is subsequently passed to the same VLA, now serving as the action policy $\pi_{\theta}$, which outputs the final robot actions. Because this spatial reasoning VQA task is specifically formulated to correlate highly with objects of interest and the requisite actions, the resulting textual pair $u^{\prime}$ provides a distilled, abstract representation of task-relevant factors, simplifies the learning challenge for the action policy, and mitigates the learning of spurious correlations.
@@ II-B2 Modeling Task-relevant Factors via Spatial Reasoning VQA
II-B3 Automated Rule-based Object Direction Labeling
To facilitate model training, ground-truth spatial relationships are integrated into the training datasets. Each training datapoint is thus augmented to $(o_{i},a_{i},u^{\prime}_{i})$, where $u^{\prime}_{i}$ encompasses self-generated questions and their corresponding ground-truth spatial relationship answers. These ground-truth relationships are derived by first determining the positions of the robot’s end-effector/gripper and relevant objects, then inferring their relative spatial arrangement, as illustrated in Fig. 3. In simulations, object locations are readily available. In real-world settings, the end-effector’s position at the instant of gripper closure/opening is recorded and used as a proxy for the object’s position when establishing these relationships. Given an end-effector position $x_{i},y_{i},z_{i}]$ and an object position $x_{0},y_{0},z_{0}]$, the position difference is calculated as $\bm{d}=[x_{i}-x_{0},y_{i}-y_{0},z_{i}-z_{0}]$. The coarse-grained spatial relationship is then determined by the axis corresponding to the component of $\bm{d}$ with the largest absolute value. For instance, if $|x_{i}-x_{0}|$ is the maximum component magnitude in $\bm{d}$, the relationship is classified as “left” or “right” based on the sign of $x_{i}-x_{0}$. When the object is grasped, as indicated by gripper closure, it is directly labeled with the text “grasped”. This rule-based strategy, applied from a third-person viewpoint, generates the ground-truth labels used for supervising the VLA’s spatial reasoning. During training, the autoregressive loss is also applied to tokens corresponding to spatial reasoning answers, alongside the primary action tokens; this additional loss encourages the model to predict the correct textual descriptions of these spatial relationships. Our experiments show that this relatively coarse-grained spatial reasoning is nonetheless effective for guiding the action generation process (see Section III-D).
III Experiments
In this section, we conduct extensive experiments to answer the following questions: 1) Can InSpire enhance VLAs on tasks from both simulation and real-world environments?
@@ III Experiments
We answer the first question in Sections III-A and III-B, the second question in Section III-C, the third question in Section III-D, and the fourth question in Section III-E.
III-A Simulation Experiments
We perform simulation experiments using LIBERO [22] and CALVIN [27] benchmarks. We train the model with random 3 seeds, and conduct 100 trials per task at inference.
III-A1 LIBERO Evaluation
We use manipulation tasks from the five datasets—LIBERO-90, LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long/10. The VLA model is jointly trained on the 90 tasks in the LIBERO-90 dataset, with each task consisting of 50 demonstrations. To comprehensively assess InSpire’s capabilities, we evaluate the model on both seen tasks from LIBERO-90 and unseen tasks from the other four datasets. We perform evaluations using two state-of-the-art models: miniVLA-VQ[3] and $\pi_{0}$-FAST [32]. In our experiments, miniVLA-VQ is pretrained from scratch, whereas $\pi_{0}$-FAST undergoes full-finetuning on the LIBERO-90 dataset. To ensure fair comparisons, both models adopt the same hyperparameters (e.g., learning rate = 1e-5 and training step = 50000) as those employed for InSpire. Since the LIBERO-90 pretrained OpenVLA model has not been released, we borrow the reported results on seen tasks from the project homepage.
TABLE I: LIBERO Performance. models borrowed from [3]. our models pretrained from scratch on LIBERO-90. our models finetuned on LIBERO-90.